find the solution x+6y=23 x-y=-5
Answer by Maths68(1474)

x+6y=23..........(1)
x-y=-5...........(2)
Take (2)
x-y=-5
x=y-5............(3)
Put the value of x from (3) to (1)
(y-5)+6y=23
y-5+6y=23
7y=23+5
7y=28
Divide by 7 both sides
7y/7=28/7
y=4
Put the value of y in (3)
x=y-5............(3)
x=4-5
x=-1
x,y=-1,4

Check
======
Put the values of x and y in (1)
x+6y=23..........(1)
-1+6(4)=23
-1+24=23
23=23
Put the values of x and y in (1)
x-y=-5...........(2)
-1-4=-5
-5=-5